OpenStudy (anonymous):

Maria is comparing the prices of two window cleaning companies. Company A charges $6 per window and an additional $12 as service charges. Company B charges $5 per window and an additional $15 as service charges. Part A: Write equations to represent Company A and Company B's total charges for cleaning a certain number of windows. Define the variable used in the equations. Part B: Which company would charge less for cleaning 8 windows? Justify your answer. Part C: How much money is saved by using the services of Company B instead of Company A to clean 6 windows?

OpenStudy (texaschic101):

Company A : 12 + 6w = total charge (w stands for number of windows) Company B : 15 + 5w = total charge (w stands for number of windows) for 8 windows : Company A : 12 + 6(8) = 12 + 48 = 60 Company B : 15 + 5(8) = 15 + 40 = 55 Company A charges less for 8 windows. for 6 windows : Company A : 12 + 6(6) = 12 + 36 = 48 Company B : 15 + 5(6) = 15 + 30 = 45 Company B charges $3 less for cleaning 6 windows

OpenStudy (texaschic101):

oops...messed up on second one....Company B charges less for 8 windows